A researcher examined health records on a sample population including 8000 women. At the end of 2010, 990 women had been diagnosed with breast cancer. During 2011, there were 10 new breast cancer cases in this population. What is the cumulative incidence for breast cancer among this population in 2011 (HINT: the number of persons at risk should be those who are breast cancer free at the beginning of 2011)? (Round to four decimal places and answer in 0.0000 format)
Answer #1

No. of women at risk at the start of 2011 is 8000-990 = 8010.

No. of new cases in 2011 = 10.

Hence cumulative incidence = 10/8010 = 0.0012 (rounded to 4 decimal places)

Cumulative incidence is the ratio of number of new cases during the period under consideration with number of person at risk in the beginning of the period under consideration.